Solve the following equation: square root of (3)*x-x=1
Answer by Theo(13342)

i believe you are saying sqrt(3) * x - x = 1
if so, then factor out the x to get (sqrt(3) - 1) * x = 1
divide both sides of the equation by (sqrt(3) - 1) to get x = (1/sqrt(3)-1))
that's your solution.
you can confirm by replacing x with 1 / (sqrt(3) - 1) in your original equation and showing that it is true.
you will get:
you can also confirm using your calculator.
take 1/(sqrt(3)-1) and get its value using the calculator.
the value will be 1.366025404
store that in your calculator memory and use it to evaluate the original equation.
i stored it into a variable called A in my calculator
that means that every time i use A, i am using 1.366025404.
the original equation is sqrt(3) * x - x = 1
using my calculator, i evaluated sqrt(3) * A - A = 1
the answer i got was 1 = 1, confirming the value of x = 1/(sqrt(3)-1) = 1.366025404 is good.
your solution is that x = 1/(sqrt(3)-1))